The equation is:
v(t)=A(1- e^(-t/t_maxspeed ) )
The tasks are to:
● Use thinking methods to analyse the given engineering problem, e.g. break the problem down into a series of manageable elements, and produce a specification
1
Expert's answer
2020-12-29T17:54:26-0500

"v(t)=A(1- e^{-t/t_{maxspeed} } )\\"


If "\pmb{t\rightarrow \infty, v(t)\rightarrow A}"

As, "\lim_{t \rightarrow \infty} v(t)=\lim_{t \rightarrow \infty} A(1-e^{-t/t_{maxspeed}})"

"=A-\lim_{t \rightarrow \infty} e^{-t/t_{maxspeed}}=A-0=A"


For "t=0, v(t)=A-A=0"

"\therefore" It starts with zero velocity.


Now, acceleration "a(t)=\frac{dv(t)}{dt}=\frac{A}{t_{maxspeed}}e^{-t/t_{maxspeed} }"


If "\pmb{t\rightarrow \infty, a(t)\rightarrow 0}"

As, "\lim_{t\rightarrow \infty}a(t)=\lim_{t\rightarrow \infty}\frac{A}{t_{maxspeed}}e^{-t/t_{maxspeed} }"

"=\frac{A}{t_{maxspeed}}\times0=0"


Now distance travelled at a time "t,"

"d(t)=\int_{0}^{t}v(t)dt=\int_{0}^{t}A(1- e^{-t/t_{maxspeed} } )dt"

"=At+At_{maxspeed}[e^{-t\/t_{maxspeed}}-1]"